OpenStudy (anonymous):

need help!! a pizza with a diameter of 18 inches is cut into 10 equal slices. To the nearest square inch, what is the area of each slice of the pizza.

OpenStudy (anonymous):

first, find the area of the circle

OpenStudy (anonymous):

So the area of each slice is \[S = \frac{ \pi•9^2 }{10} or \frac{ \pi•81 }{ 10 }\]

OpenStudy (anonymous):

You can figure this out on paper or with calculator. Substitute 3.14 as pi. It will make it easier

OpenStudy (anonymous):

@briannemarie